1. You are the manager of the Sea Breeze Motel and one night you rented 5 singles and 12 doubles for a total of $1515. The next night you rented 9 singles and 10 doubles for a total of $1625. What is the rental charge for each type of room? You must identify each equation and solve using either substitution or elimination.
Answer by checkley77(12844)

5S+12D=1,515 MULTIPLY BY 9
9S+10D=1,625 MULTIPLY BY -5 & SUBTRACT.
45S+108D=13,635
-45S-50D=-8,125
--------------------
58D=5,510
D=5,510/58
D=$95 PRICE OF A DOUBLE ROOM.
5S+12*95=1,515
5S+1,140=1,515
5S=1,515-1,140
5S=375
S=375/5
S=$75 PRICE OF A SINGLE ROOM.
PROOF:
9*75+10*95=1,625
675+950=1,625
1.625=1,625